OpenStudy (anonymous):

Exponential equation. Points (-2,9) (0,1). My answer y=3(1/3)^x. Am I correct?

OpenStudy (anonymous):

Well \[3*(1/3)^{-2} = 3*3^2 = 9*3 = 27 \ne 9\] so nope. That can't be right.

OpenStudy (anonymous):

hmm well lets see. First I got a=9/b^2. is that correct?

OpenStudy (anonymous):

From the second point we have \[1 = ab^0\] Since \[b^0=1\rightarrow 1 = ab^0 = a*1 \rightarrow a = 1\]

OpenStudy (anonymous):

Try plotting y = 2^x, y = 2^(-x), etc. That may give you ideas about the base and the exponent,

OpenStudy (anonymous):

oh well I was taught to use the first point first to find out what a equals for the b equation.

OpenStudy (anonymous):

....? plotting?

OpenStudy (anonymous):

Anytime you have x = 0 for an exponential equation, the y value for that point is the value of a.

OpenStudy (anonymous):

Because any base raised to the 0 power equals 1.

OpenStudy (anonymous):

still no comprendo senor.....or senora. You don't use the second equation for a.......gawh I wish math was simple for fooligans like me.

OpenStudy (anonymous):

Ok. Lets break it down. We have 2 points (-2,9) and (0,1) That gives 2 equations: \[9 = ab^{-2}\] and \[1 = ab^0\]

OpenStudy (anonymous):

With me so far?

OpenStudy (anonymous):

si.

OpenStudy (anonymous):

Ok. So what is \[b^0=?\]

OpenStudy (anonymous):

this is where your throwing me off. Don't i have to use the first equation for a first?

OpenStudy (anonymous):

It doesn't matter. You use whichever is easier. In this case the second equation is easier. This will become apparent when you understand what b^0 is.

OpenStudy (anonymous):

okay. I just want to make sure. I have a test tomorrow and if i use your tricks I don't want to get penalized.

OpenStudy (anonymous):

Nope. Not a trick.

OpenStudy (anonymous):

okay well to me its a different way so continue senor.

OpenStudy (anonymous):

If you really want we can do it your way.

OpenStudy (anonymous):

no I'm open to your idea. your the smarter one here so continue.

OpenStudy (anonymous):

Lets do it your way and then we'll do mine and see that we get the same thing.

OpenStudy (anonymous):

okay sounds good to me. :0]

OpenStudy (anonymous):

Are you familiar with this notation? \[\rightarrow\]

OpenStudy (anonymous):

does that mean equals or references to the next step?

OpenStudy (anonymous):

not equals. It means 'implies', 'therefore', or 'leads to'

OpenStudy (anonymous):

So something like \[ A = 5\] \[ A +B = 10 \rightarrow 5 + B = 10 \rightarrow B = 5\]

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

okay continue :0]

OpenStudy (anonymous):

So using the first equation and solving for a: \[9 = ab^{-2}\] \[\rightarrow 9b^2 = ab^{-2}b^2\] \[\rightarrow9b^2 = ab^0 \rightarrow a = 9b^2\]

OpenStudy (anonymous):

Follow that?

OpenStudy (anonymous):

Okay question time! Don't you divide by b^2? and keep 9/b^2 for the next equations a? for example: 1=(9/b^2)(b^0)

OpenStudy (anonymous):

then solve for b and then plug in b for the original equation.

OpenStudy (anonymous):

Ah. No because \[b^{-2} = \frac{1}{b^2}\]

OpenStudy (anonymous):

So you need to multiply by b^2 in order to get a by itself, not divide.

OpenStudy (anonymous):

oooooh so thats where I missed the mark. Okay i get it so if i have a negative, instead of dividing I multiply.

OpenStudy (anonymous):

Think about it this way. You want to move the b over to the other side. You know that multiplying powers of the same base you add exponents right? \[a^b * a^c = a^{(b+c)}\]

OpenStudy (anonymous):

yes and when you divide you subtract.

OpenStudy (anonymous):

Right. So here's the million dollar question. What is \[k^0\]

OpenStudy (anonymous):

0 right?

OpenStudy (anonymous):

Nope. 2^4 = 16 2^3 = 8 2^2 = 4 2^1 = ? 2^0 = ?

OpenStudy (anonymous):

shoot I change my mind its 1!! now wheres my money?

OpenStudy (anonymous):

heheh.

OpenStudy (anonymous):

Right. So if we have \[ab^2 = 9\] We can multiply both sides by b^-2 \[a*b^2*b^{-2} = 9*b^{-2}\]

OpenStudy (anonymous):

And we end up with \[a*b^{2+ (-2)} = 9b^{-2} \] Which simplifies to \[a*b^{0} = 9b^{-2} \] Which simplifies to \[a*1 = 9b^{-2} \] Which simplifies to \[a = 9b^{-2} \]

OpenStudy (anonymous):

Now that wasn't the equation we had. We had \[ab^{-2} = 9\] So instead of multiplying by b^-2 we need to multiply by b^2

OpenStudy (anonymous):

okay gotcha.

OpenStudy (anonymous):

So we end up with \[a = 9b^2\]

OpenStudy (anonymous):

And we plug that into the second equation.

OpenStudy (anonymous):

So what do you get when you plug in 9b^2 for a in the second equation.

OpenStudy (anonymous):

i got 1/9=b^2

OpenStudy (anonymous):

after dividing 9 from 9b^2 of course.

OpenStudy (anonymous):

So b is?

OpenStudy (anonymous):

b=1/3 Like I had before...

OpenStudy (anonymous):

Yeah. So now plug that in for b in either of the two equations.

OpenStudy (anonymous):

then i got a=1. and a full answer of y=1(1/3)^x

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

So now look at what happens when you plug in 1/3 for b in the second equation. Does it matter what b was?

OpenStudy (anonymous):

what do you mean? I got the whole equation right didnt i?

OpenStudy (anonymous):

Yes. I'm trying to show my original idea.

OpenStudy (anonymous):

oh okay. well no it doesn't matter what b was. So yes you were right. hahaa that all you wanted to hear wasnt it?

OpenStudy (anonymous):

\[1=ab^0\] is our second equation from the beginning. You plug in 1/3 and you get a = 1. But does it matter what you plug in for b?

OpenStudy (anonymous):

no just as long as the answer is correct. :0]

OpenStudy (anonymous):

Right. So you can see that for exponential equations if you are given a point (0,k). You already know that a must be k.

OpenStudy (anonymous):

oh okay yeah I get it.

OpenStudy (anonymous):

Because \[ab^0= k\rightarrow a*1 = k \rightarrow a = k\]

OpenStudy (anonymous):

Which means you can jump right to finding b and save some time.

OpenStudy (anonymous):

so if i do it either way i will end up with the right answer. Okay i'm getting your drift here. thanks a bunch! your so patient and I feel more confident it passing my test. SO thanks it really means a lot!!

OpenStudy (anonymous):

Anytime =)